OpenStudy (anonymous):

15=5^(x+3)-7 please help

OpenStudy (jdoe0001):

subtract 7 from both sides

OpenStudy (jdoe0001):

well, add rather :/

OpenStudy (anonymous):

then

OpenStudy (jdoe0001):

\(\bf 15=5^{x+3}-7 \implies 22 = 5^{x+3}\\ \textit{log cancellation rule of } log_aa^x = x\\ log_5{22} = log_5(5^{x+3}) \implies \square? = \square?\)

OpenStudy (jdoe0001):

or lemme put it this way \(\bf log_\color{red}{a}\color{red}{a}^x = x \implies log_\color{red}{5}(\color{red}{5}^{x+3}) = \square?\)